THE ν_{16} BAND IN TRANS 1,2 DICHLOROETHANE

Y. EL YOUSSOUFI, \underline{R. GEORGES}, J. LIEVIN, J. VANDER AUWERA and M. HERMAN, Laboratoire de Chimie Physique Moléculaire, CP160/09, Université Libre de Bruxelles, B-1050, Belgium.

The ν_{16} band in trans 1,2 dichloroethane at 1232 cm^{−1} was recorded at high resolution using a Fourier transform interferometer. Room temperature and jet-cooled experimental conditions were used. The experimental set-up allowing the jet-cooled spectra to be recorded will be briefly presented. The comparison between both sets of data leads to identify fundamental and hot bands for various isotopomers. Assignments are suggested for most of them on the basis of a detailed investigation of the Q–branch contours. Anharmonicity parameters are obtained. The rotational analysis of the fundamental band in two isotopomers is performed using the room temperature data.